Circular No. 8356

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 L2 (LINEAR) An apparently asteroidal object reported by LINEAR (discovery ob-

servation below), and posted on the ‘NEO Confirmation Page’, has been found to show cometary appearance on CCD images taken by G. Masi (Las Campanas, 0.35-m reflector, June 17.2 UT; round coma with diameter 6′′- 8′′) and by J. E. McGaha (Tucson, AZ, 0.62-m reflector, June 17.3; 5-min exposure shows a 10′′ coma and a 15′′ fan-shaped tail in p.a. 300^{o}).

2004 UT  α_{2000}  δ_{2000}  Mag.
June 11.35259  19^{h}10^{m}44.^{s}61  −27^{o}36′24′′.0  19.2

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2004-M09.

T = 2005 Nov. 16.015 TT  ω = 256.^{o}956
Ω = 99.245 }2000.0
q = 3.80396 AU  i = 62.969

COMETS C/2003 V8, W3–W8, X1–X4 (SOHO) Further to IAUC 8354, K. Battams reports measurements for addi-

tional Kreutz sungrazing comets found on SOHO website C2 images by J. Sachs (C/2003 V8, W6), X.-m. Zhou (C/2003 W3, W7, X3), D. Evans (C/2003 W4), M. Boschat (C/2003 W5, X2), G. Gallina (C/2003 W8), T. Hoffman (C/2003 X1), and R. Kracht (C/2003 X4). C/2003 V8 and W6 were also visible on C3 images. The reductions and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  2003 UT  α_{2000}  δ_{2000}  MPEC
C/2003 V8 Nov. 13.471  14^{h}57.^{m}2  −21^{o}23′  2004-K54
C/2003 W3  16.561  15 23.4  −20 12  2004-K54
C/2003 W4  21.743  15 45.6  −21 32  2004-K54
C/2003 W5  22.754  15 51.7  −21 54  2004-K54
C/2003 W6  25.571  15 49.1  −26 06  2004-K55
C/2003 W7  27.326  16 10.9  −22 47  2004-K55
C/2003 W8  29.435  16 19.3  −22 58  2004-K55
C/2003 X1 Dec. 1.160  16 28.1  −23 26  2004-K55
C/2003 X2  1.561  16 32.4  −23 38  2004-K55
C/2003 X3  2.201  16 33.1  −23 39  2004-K55
C/2003 X4  4.601  16 43.7  −24 03  2004-L24
2004 June 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green